OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 6 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (nurali):

V = 3.14r^2h /3 V = 3.14*6^2*h / 3 = 300 km^3 Multiply both sides by 3,solve for H 3.14*36*H = 900 H = 900 / (3.14*36) = 7.96 km^3

OpenStudy (anonymous):

Thank You

OpenStudy (nurali):

Anytime.